OpenStudy (anonymous):

What is the value of the function y = 0.5x – 10 when x = 9? y = ___ @Legends @Nibby

OpenStudy (anonymous):

For this problem, we're just going to substitute 9 for x, so the equation will look like:\[y = 0.5(9) - 10\] To solve this, simply times 0.5 by 9, and then minus the 10 from the number to get your Y. Can you solve this?

OpenStudy (anonymous):

yes! thanks! @Nibby

OpenStudy (anonymous):

You're welcome :) If I helped, make sure to hit 'Best Response' and close the question.

OpenStudy (anonymous):

so the answer is -5.5? @Nibby

OpenStudy (anonymous):

Well, to find it, we do:\[0.5 * 9 = -4.5\] Next, we plug that into the equation to get: \[y = 4.5 - 10\] If we plug 4.5 - 10 into a calculator, we get -5.5. So, yes! Your answer is -5.5. Good job!